# Enhancing Zero-Shot Facial Expression Recognition by LLM Knowledge Transfer

### v2

@@ Abstract
Current facial expression recognition (FER) models are often designed in a supervised learning manner and thus are constrained by the lack of large-scale facial expression images with high-quality annotations. Consequently, these models often fail to generalize well, performing poorly on unseen images in inference. Vision-language-based zero-shot models demonstrate a promising potential for addressing such challenges. However, these models lack task-specific knowledge and therefore are not optimized for the nuances of recognizing facial expressions. To bridge this gap, this work proposes a novel method, Exp-CLIP, to enhance zero-shot FER by transferring the task knowledge from large language models (LLMs). Specifically, based on the pre-trained vision-language encoders, we incorporate a projection head designed to map the initial joint vision-language space into a space that captures representations of facial actions. To train this projection head for subsequent zero-shot predictions, we propose to align the projected visual representations with task-specific semantic meanings derived from the LLM encoder, and the text instruction-based strategy is employed to customize the LLM knowledge. Given unlabelled facial data and efficient training of the projection head, Exp-CLIP achieves superior zero-shot results to the CLIP models and several other large vision-language models (LVLMs) on seven in-the-wild FER datasets. The code and pre-trained models are available at https://github.com/zengqunzhao/Exp-CLIP.
Index Terms:
Facial expression recognition, zero-shot learning, large language models, contrastive learning
I Introduction
@@
Facial expressions are a crucial aspect of human communication, allowing individuals to convey a wide range of emotions and intentions non-verbally. Facial expression recognition (FER) refers to the identification and analysis of human facial expressions to understand the underlying emotions or sentiments, which has various applications such as enabling people who are blind or have low vision to understand the emotions of people around them better [1], monitoring patient emotions for better mental health assessment [2] and improving user experience in applications and devices [3]. FER can be categorized into two types based on input distinctions: Static FER (SFER) and Dynamic FER (DFER). SFER involves analyzing individual images or snapshots of faces while DFER focuses on analyzing facial expressions across a sequence of images or frames.
In the past decade, research in both SFER and DFER has shifted from controlled laboratory environments to more challenging in-the-wild conditions [4]. Given the availability of large-scale facial expression datasets captured in real-world scenarios [5, 6, 7, 7] and advancements in foundational models [8, 9], current FER models have successfully attained satisfactory Intradomain performance [10, 11, 12, 13]. However, these deep FER models were mainly trained in a supervised manner, relying heavily on large-scale training data with categorical annotations. To make further improvements, extensive labelled training data are increasingly demanded, which is both labour-intensive and expensive. Furthermore, these supervised FER models often fail to generalize well from one domain to another [14]. Additionally, FER annotations frequently encounter challenges related to noise and inconsistent annotations caused by subjectivity, hindering the training process, and different datasets may use different affect models [15, 16, 17, 18], making the issue more challenging. In contrast, zero-shot learning uses semantic information to bridge the gap between the seen and unseen classes, is capable of generalization to unseen expressions and removes the need to annotate the training data necessary in the supervised learning paradigm [19].
@@
Recently, large vision-language pre-training models have shown potential for achieving promising zero-shot performance [20, 21, 22, 23]. Those models are often designed using one of the following two strategies: (a) aligning large connections of images and raw texts using two separate encoders [20, 21, 24] or (b) incorporating visual modality to Large Language Models (LLMs) to create Large Vision-Language Models (LVLMs) [23, 25, 26]. Under the former training paradigm, the CLIP model [20] was proposed to capture the relationship between large-scale images and their associated textual descriptions, achieving outstanding zero-shot performance on various classification tasks. Building upon the CLIP model, the FaRL model [21] is learned from face-image-text pairs, producing a universal facial representation, which shows superior results on downstream face analysis tasks. EmotionCLIP [27] is proposed to extract visual emotion representations from verbal and nonverbal communication. However, the original CLIP model learned more general feature representations may not be optimized for the nuances of recognizing facial expressions (see Fig. 1). Training texts in FaRL and EmotionCLIP are either general appearance descriptions or conversations, lacking detailed facial action of expression descriptions. Fine-tuning the CLIP model with facial expression visual data and their corresponding text descriptions, as discussed in [28], appears to be a more direct approach. However, this idea relies on facial images with detailed facial action descriptions and has the same manual labour costs as the supervised approaches. Whilst LVLM-based approaches appear to promise automation, prompting such models for zero-shot tasks demands substantial engineering efforts [29]. Additionally, these models suffer problems caused by hallucinations of LLM decoders[30]. It remains a significant challenge to achieve reliable zero-shot FER.
@@
  • The proposed method outperforms the various large vision-language models (LVLMs) on both image-based and video-based FER datasets.
II Related Work
II-A Facial Expression Recognition
@@
Traditional FER Models: The traditional FER model relies on a feature extractor and a classifier. In the past decade, research on static FER has primarily focused on addressing real-world challenges such as occlusion [31], pose variation [32], domain shift [14], and label noise [33, 34, 35]. In the latest work, Wu et al.[35] introduce a landmark-aware net to address the ambiguity of expressions by leveraging facial landmarks to mitigate the impact of label noise. Lee et al.[36] proposed LatentOFER to improve FER accuracy under face occlusions. In their method, the model can detect occlusions, restore occluded parts of the face as if they were unoccluded, and recognize them. To address the imbalanced FER problem, Zhang et al.[37] propose to leverage re-balanced attention maps to regularize the model, enabling it to extract transformation invariant information about the minor classes from all training samples. In contrast, research on Dynamic FER primarily aims to develop models capable of generating robust spatial-temporal representations of facial expressions. The introduction of the Transformer to DFER, as seen in Former-DFER [38], marked a significant advancement, demonstrating superior performance compared to methods based on 3DCNNs or CNN-RNNs. Subsequent studies have further enhanced performance by addressing issues such as noisy frames [39] and the intensity of varied frames [12]. Moreover, Wang et al.[13] introduced M3DFEL to tackle the challenge of imbalances between short- and long-term temporal relationships in DFER. The concept of soft labelling [40] has also been applied to reduce data ambiguity.
@@
Vision-language-based FER Models: In contrast to traditional FER models, vision-language-based models do not rely on classifiers. Instead, they derive final classification results by comparing the similarity between visual features and diverse text features, selecting the highest similarity as the ultimate prediction. Li et al.[41] introduced CLIPER, a unified framework for both static and dynamic FER based on CLIP. CLIPER incorporates multiple expression text descriptors to learn fine-grained expression representations, achieving state-of-the-art performance on various FER benchmarks. Zhao et al.[42] proposed DFER-CLIP, designed for dynamic FER with temporal learning, introducing textual descriptions of facial behaviour for contrastive learning. Zhang et al.[11] presented the CLEF model, which conducts text-driven contrastive learning for facial behaviour understanding, utilizing activity descriptions and coarse-grained information from certain datasets. Nevertheless, the reliance on extensive labelled facial data in these models presents a challenge in data acquisition. The recent work by Foteinopoulou et al.[28] introduces a novel vision-language model utilizing sample-level text descriptions as natural language supervision for facial expression video to enhance zero-shot FER classification. Nevertheless, this approach faces the difficulty of labelling detailed facial action descriptions, requiring expertise in perceiving nuanced facial movements. In contrast, our method utilizes facial images without annotations in an unsupervised fashion, presenting a more labour-efficient approach.
II-B Large Vision-Language Models
@@
The current landscape of LVLMs can be broadly categorized into two main types. The first type involves aligning large collections of images and raw texts by utilizing two distinct encoders [20, 43, 44, 21, 24, 27, 45]. The second type integrates the visual modality into LLMs to form comprehensive LVLMs [26, 23, 25]. In the context of the first type, the pioneer work CLIP [20] maps data of different modalities, text and images, into a shared embedding space. This shared space exhibits the capability to generalize to various image classification tasks with zero- and few-shot learning. Subsequently, VideoCLIP [43] and XCLIP [44] were introduced for zero-shot video-text understanding. FaRL [21] focuses on learning universal facial representations from face-image-text pairs, while EmotionCLIP [27] is designed to extract visual emotion representations from both verbal and nonverbal communication. The recent FLAVA [24] learned a foundational language and vision representation that enables unimodal vision and language understanding as well as multimodal reasoning, all within a single pre-trained model. Regarding the LLMs-based models, Flamingo [26] bridges the powerful pre-trained vision-only and language-only models, and BLIP2 [23] further improves the capacity of the LVLMs by bridging the modality gap with a lightweight Querying Transformer and adopting more powerful LLM. The recent model LLaVa [25] was trained by fine-tuning LLMs on GPT-generated multimodal instruction-following data, demonstrating impressive multimodel chat abilities. A recent work [46] also leverages the common sense knowledge and reasoning ability of the LLaVa and shows that this knowledge can be transferred to the emotion recognition problem. Nevertheless, the zero-shot FER performance of these models is challenged by substantial engineering efforts on prompting [29] and the hallucination issue inherited from LLMs decoders[30].
III Method
@@
The core idea of our method is to map a general visual representation into a task-aware representation beneficial to facial expression recognition. A projection-based strategy is proposed to fulfil this purpose, in which a projection head is learned through leveraging the task-specific knowledge from the LLMs. We will first present an overview of our method in section III-A and then illustrate the zero-shot FER task in section III-B, finally, we will show how to train the learnable projection head for zero-shot learning in section III-C.
@@
III-A Overview
III-B Enhanced Zero-shot FER
@@
III-C LLM Knowledge Transferring
The plain CLIP feature space learns general representations, lacking task-specific knowledge [44, 21]. Therefore, we leverage the task-specific knowledge generated from LLMs to improve the CLIP feature representations. This task-specific knowledge is elicited by an instruction-based strategy.
@@ III-C LLM Knowledge Transferring
IV Experiments
IV-A Datasets
To verify the effectiveness of the proposed method, we conduct experiments on seven popular in-the-wild facial expression benchmarks, including three static FER datasets: RAF-DB [5], AffectNet [6], and FERPlus [47], and four dynamic FER datasets: DFEW [7], FERV39k [48], MAFW [49], and AFEW [50].
@@ IV-A Datasets
IV-B Implementational Details
Experiment Settings: In the contrastive learning phase, the training is conducted on CAER-S [51] training set with 44,996 facial images. The parameters are optimized by the SGD optimizer with a mini-batch size of 512. The projection head in our method refers to a matrix with size $512\times 4096$ when utilising ViT-B-32/B-16 or size $768\times 4096$ when utilising ViT-L-14. The learning rate for the learnable projection head is set as $1\times 10^{-3}$. The learnable projection head is trained for 5 epochs in an end-to-end manner. Our models are trained on the Tesla A100 GPU based on the open-source PyTorch platform. We employ the Flan-T5-XXL [52] with 11B parameters as our LLM. During the zero-shot prediction, for the SFER, all the images are resized to $224\times 224$ as an input. For the DEER, all the fixed 16 frames of sequence in our experiments follow the sampling strategy in [7, 38, 12], and then resized to $224\times 224$ as inputs. To obtain more stable and reliable results, we train models three times with different random seeds and then use the average as the final result.
IV-C Comparisons with the State-of-the-Art
TABLE I: Zero-shot FER results in comparison with other visual language models. B/32, B/16, L/14 denote ViT-B/32, ViT-B/16, and ViT-L/14, respectively. EmoCLIP was pre-trained on MAFW [49] with facial expression clips and action unit text descriptions without categorical labels. Bold donates the best, Underline denotes the second best.
@@ IV-C Comparisons with the State-of-the-Art
TABLE II: Zero-shot FER results in comparison with CLIP model. All models use the prompt “a photo of a face with an expression of [class].” for zero-shot prediction.

### v3

@@ Abstract
Current facial expression recognition (FER) models are often designed in a supervised learning manner and thus are constrained by the lack of large-scale facial expression images with high-quality annotations. Consequently, these models often fail to generalize well, performing poorly on unseen images in inference. Vision-language-based zero-shot models demonstrate a promising potential for addressing such challenges. However, these models lack task-specific knowledge and therefore are not optimized for the nuances of recognizing facial expressions. To bridge this gap, this work proposes a novel method, Exp-CLIP, to enhance zero-shot FER by transferring the task knowledge from large language models (LLMs). Specifically, based on the pre-trained vision-language encoders, we incorporate a projection head designed to map the initial joint vision-language space into a space that captures representations of facial actions. To train this projection head for subsequent zero-shot predictions, we propose to align the projected visual representations with task-specific semantic meanings derived from the LLM encoder, and the text instruction-based strategy is employed to customize the LLM knowledge. Given unlabelled facial data and efficient training of the projection head, Exp-CLIP achieves superior zero-shot results to the CLIP models and several other large vision-language models (LVLMs) on seven in-the-wild FER datasets. The code is available at https://github.com/zengqunzhao/Exp-CLIP.
1 Introduction
@@
Facial expressions are crucial for human communication, allowing individuals to convey a wide range of information non-verbally. In the field of Human Computer Interaction, Facial Expression Recognition (FER) often aims at the analysis of the underlying emotions or sentiments, and has various applications such as assisting people who are blind or have low vision [12], monitoring patients for better mental health assessment [3] and improving user experience when interacting with applications and devices [24]. FER can be categorized into two types based on input distinctions: Static FER (SFER) and Dynamic FER (DFER). SFER involves analyzing individual images or snapshots of faces while DFER focuses on analyzing facial expressions across a sequence of images or frames.
@@
In the past decade, research in both SFER and DFER has shifted from controlled laboratory environments to more challenging in-the-wild conditions [22]. Given the availability of large-scale facial expression datasets captured in real-world scenarios [21, 25, 11, 11] and advancements in foundational models [10, 7], current FER models have attained high accuracy [44, 42, 17, 33]. However, these deep FER models were mainly trained in a supervised manner, relying heavily on large-scale training data with categorical annotations. The reliance on such labelled data presents significant challenges. Firstly, the process of acquiring and annotating large-scale training data is both labour-intensive and costly. Secondly, supervised FER models tend to struggle with generalization across different domains [4]. This means that while these models may perform well within the domain they were trained on, their performance often degrades when applied to new, unseen domains. Additionally, FER annotations frequently encounter challenges related to noise and inconsistent annotations caused by subjectivity, hindering the training process [34, 38]. Given these challenges, exploring alternative approaches becomes crucial. One promising direction is zero-shot learning, which uses semantic information to bridge the gap between seen and unseen classes. This approach is capable of generalizing to unseen expressions and removes the need for extensive annotation of training data [27].
Recently, large vision-language pre-training models have shown potential for achieving promising zero-shot performance [28, 49, 37, 20]. Those models are often designed using one of the following two strategies: (a) aligning large connections of images and raw texts using two separate encoders [28, 49, 30] or (b) incorporating visual modality to Large Language Models (LLMs) to create Large Vision-Language Models (LVLMs) [20, 16, 1]. Under the previous training paradigm, the CLIP model [28] was designed to learn from vast datasets containing millions of image-text pairs. This extensive training enables the model to achieve remarkable zero-shot capabilities, effectively categorizing images into various classes without specific training on those categories. Building on the foundations of the CLIP model, the FaRL model [49] extends this approach by focusing specifically on face-image-text pairs. This adaptation allows FaRL to develop a universal facial representation effective across a broad spectrum of face analysis tasks. Compared to previous models, FaRL demonstrates enhanced performance in accuracy and generalization. EmotionCLIP [41] is proposed to extract visual emotion representations from verbal and nonverbal communication. However, the original CLIP model learned more general feature representations may not be optimized for the nuances of recognizing facial expressions (see Fig. 1). Training texts in FaRL and EmotionCLIP are either general appearance descriptions or conversations, lacking detailed facial action of expression descriptions. Fine-tuning the CLIP model with facial expression visual data and their corresponding text descriptions, as discussed in [8], appears to be a more direct approach. However, this idea relies on facial images with detailed facial action descriptions and has the same manual labour costs as the supervised approaches. Whilst LVLM-based approaches appear to promise automation, prompting such models for zero-shot tasks demands substantial engineering efforts [9]. Additionally, these models suffer problems caused by hallucinations of LLM decoders[29]. It remains a significant challenge to achieve reliable zero-shot FER.
@@
  • The proposed method outperforms the various LVLMs on both image-based and video-based FER datasets.
2 Related Work
2.1 Facial Expression Recognition
@@
Traditional FER Models: The traditional FER model relies on a feature extractor and a classifier. In the past decade, research on static FER has primarily focused on addressing real-world challenges such as occlusion [35, 46], pose variation [32], domain shift [4], and label noise [34, 38, 47]. In the latest work, Wu et al. [38] introduce a landmark-aware net to address the ambiguity of expressions by leveraging facial landmarks to mitigate the impact of label noise. Lee et al. [14] proposed LatentOFER to improve FER accuracy under face occlusions. In their method, the model can detect occlusions, restore occluded parts of the face as if they were unoccluded, and recognize them. To address the imbalanced FER problem, Zhang et al. [43] propose to leverage re-balanced attention maps to regularize the model, enabling it to extract transformation invariant information about the minor classes from all training samples. In contrast, research on Dynamic FER primarily aims to develop models capable of generating robust spatial-temporal representations of facial expressions. The introduction of the Transformer to DFER, as seen in Former-DFER [45], marked a significant advancement, demonstrating superior performance compared to methods based on 3DCNNs or CNN-RNNs. Subsequent studies have further enhanced performance by addressing issues such as noisy frames [19] and the intensity of varied frames [17]. Moreover, Wang et al. [33] introduced M3DFEL to tackle the challenge of imbalances between short- and long-term temporal relationships in DFER. The concept of soft labelling [13] has also been applied to reduce data ambiguity.
@@
Vision-language-based FER Models: In contrast to traditional FER models, vision-language-based models do not rely on classifiers. Instead, they derive final classification results by comparing the similarity between visual features and diverse text features, selecting the highest similarity as the ultimate prediction. Based on the CLIP, Li et al. [18] introduced CLIPER, which incorporates multiple expression text descriptors to learn fine-grained expression representations, achieving state-of-the-art performance on various FER benchmarks. Zhao et al. [48] proposed DFER-CLIP, designed for dynamic FER with temporal learning, introducing textual descriptions of facial behaviour for contrastive learning. Zhang et al. [42] presented the CLEF model, which conducts text-driven contrastive learning for facial behaviour understanding, utilizing activity descriptions and coarse-grained information from certain datasets. However, the reliance on extensive labelled facial data in these models presents a challenge in data acquisition. The recent work by Foteinopoulou et al. [8] introduces a novel vision-language model utilizing sample-level text descriptions as natural language supervision for facial expression video to enhance zero-shot FER classification. Nevertheless, this approach faces the difficulty of labelling detailed facial action descriptions, requiring expertise in perceiving nuanced facial movements. In contrast, our method utilizes facial images without annotations in an unsupervised fashion, presenting a more labour-efficient approach.
2.2 Large Vision-Language Models
@@
The current landscape of LVLMs can be broadly categorized into two main types. The first type involves aligning large collections of images and raw texts by utilizing two distinct encoders [28, 40, 26, 49, 30, 41]. The second type integrates the visual modality into LLMs to form comprehensive LVLMs [1, 20, 16, 39]. In the context of the first type, the pioneer work CLIP [28] maps data of different modalities, text and images, into a shared embedding space. This shared space exhibits the capability to generalize to various image classification tasks with zero- and few-shot learning. Subsequently, VideoCLIP [40] and XCLIP [26] were introduced for zero-shot video-text understanding. FaRL [49] focuses on learning universal facial representations from face-image-text pairs, while EmotionCLIP [41] is designed to extract visual emotion representations from both verbal and nonverbal communication. The recent FLAVA [30] learned a foundational language and vision representation that enables unimodal vision and language understanding as well as multimodal reasoning, all within a single pre-trained model. Regarding the LLMs-based models, Flamingo [1] bridges the powerful pre-trained vision-only and language-only models, and BLIP2 [20] further improves the capacity of the LVLMs by bridging the modality gap with a lightweight Querying Transformer and adopting more powerful LLM. The recent model LLaVa [16] was trained by fine-tuning LLMs on GPT-generated multimodal instruction-following data, demonstrating impressive multimodel chat abilities. Nevertheless, the zero-shot FER performance of these models is challenged by substantial engineering efforts on prompting [9] and the hallucination issue inherited from LLMs decoders[29].
3 Methods
@@
The core idea of our method is to map a general visual representation into a task-aware representation beneficial to facial expression recognition. A projection-based strategy is proposed to fulfil this purpose, in which a projection head is learned through leveraging the task-specific knowledge from the LLMs. We will first present an overview of our method in Sec. 3.1 and then illustrate the zero-shot FER task in Sec. 3.2, finally, we will show how to train the learnable projection head for zero-shot learning in Sec. 3.3.
3.1 Overview
@@
3.2 Enhanced Zero-shot FER Model
@@
3.3 LLM Knowledge Transferring
The plain CLIP feature space learns general representations, lacking task-specific knowledge [26, 49]. Therefore, we leverage the task-specific knowledge generated from LLMs to improve the CLIP feature representations. This task-specific knowledge is elicited by an instruction-based strategy. To enhance the CLIP visual feature representation for understanding facial activities, we propose projecting the original CLIP visual features into a domain-aware feature space using a projection head, rather than fine-tuning the CLIP image encoder. The key design principle of this projection head is to maintain the strong vision-language alignment inherent in CLIP while emphasizing crucial task-specific elements and filtering out irrelevant information. To achieve this, we employ a linear layer as the projection head, implemented using a simple MLP layer. The empirical results in 4.3 demonstrate that this single MLP layer outperforms more complex modules, providing an efficient and effective way to improve zero-shot performance.
@@ 3.3 LLM Knowledge Transferring
4 Experiments
4.1 Datasets
We conduct experiments on seven popular in-the-wild facial expression benchmarks, including three static FER datasets: RAF-DB [21], AffectNet [25], and FERPlus [2], and four dynamic FER datasets: DFEW [11], FERV39k [36], MAFW [23], and AFEW [6].
@@ 4.1 Datasets
Table 1: Zero-shot FER results in comparison with other visual language models. B/32, B/16, L/14 denote ViT-B/32, ViT-B/16, and ViT-L/14, respectively. EmoCLIP was pre-trained on MAFW [23] with facial expression clips and action unit text descriptions without categorical labels. Bold donates the best, Underline denotes the second best.
@@
Table 2: Zero-shot FER results in comparison with CLIP model.
@@
Experiment Settings: In the contrastive learning phase, the training is conducted on CAER-S [15] training set with 45K facial images. The projection head in our method refers to a matrix with size $512\times 4096$ when utilising ViT-B-32/B-16 or size $768\times 4096$ when utilising ViT-L-14. The learning rate for the learnable projection head is set as $1\times 10^{-3}$. The learnable projection head is trained for 5 epochs in an end-to-end manner. The parameters of the projection head are optimized by the SGD optimizer with a mini-batch size of 512. Our models are trained on the Tesla A100 GPU based on the open-source PyTorch platform. We employ the Flan-T5-XXL [5] with 11B parameters as our LLM. During the zero-shot prediction, for the SFER, all the images are resized to $224\times 224$ as an input. For the DEER, all the fixed 16 frames of sequence in our experiments follow the sampling strategy in [11, 45, 17], and then resized to $224\times 224$ as inputs. To obtain more stable and reliable results, we train models three times with different random seeds and then use the average as the final result.
4.2 Comparisons with the State-of-the-Art
BLIP2 [20] is a state-of-the-art model on various vision-language tasks, to test its zero-shot performance on the FER task, inspired by work [37], we take the facial images as input and prompt BLIP2 by using: “Question: Play the role of a facial expression classification expert and select one category with the highest similarity to the input image. Here are the optional categories: ‘contempt’, ‘fear’, ‘disgust’, ‘surprise’, ‘anger’, ‘neutral’, ‘sadness’, ‘happiness’. Answer:”. We then calculate the WAR and UAR based on the output. As shown in Tab. 1, our results outperform the BLIP2 model on all the FER test sets. Specifically, our ViT-B-16-based zero-shot results can beat BLIP2 in most SFER test sets. Furthermore, we also compare our results with several other visual-language models. VideoCLIP [40], XCLIP [26], and EmotionCLIP [41] are proposed for video analysis, video recognition, and video emotion recognition, respectively, in which VideoCLIP was pre-trained on video clips and text transcriptions, XCLIP was pre-trained on action recognition dataset with action labels, and EmotionCLIP was pre-trained on the videos and conversations. These models all learn spatial-temporal features during the training. However, our method with the bare static model can still outperform them on four DFER test sets. Moreover, we also compared our method with CLIP betterment: FaRL and FLAVA, in which the FaRL model was pre-trained on LAION-FACE specializing in the face analysis task and FLAVA improves CLIP zero-shot performance by introducing both unimodal and multimodal architecture. The results in Tab. 1 demonstrate our results are superior to these two methods on all of the FER test sets. Compared with the EmoCLIP [8], which was trained on the MAFW clips and facial action descriptions, our method outperforms it on DFEW, FERV39k, and AFEW test sets in five metrics without explicit text descriptions, indicating the effectiveness of the proposed method.
4.3 Ablation Analysis
Evaluation of the Projection Head Effectiveness: Our method is built on the CLIP model [28], to evaluate the effectiveness of the learned projection head, we compare our method to the CLIP model from two views: various ViT models and various zero-shot prompts. Our method uses the instruction “Please play the role of a facial action describer. Objectively describe the detailed facial actions of the person in the image.”. As shown in Tab. 2, under different ViT models, our method outperforms the plain CLIP model on all the test sets. With ViT-B-32, our method improves the average UAR by 2.71% and average WAR by 4.42% over eight test sets. With ViT-B-16, our method improves the average UAR by 5.47% and average WAR by 6.75% over eight test sets. With ViT-L-14, our method improves the average UAR by 8.07% and average WAR by 8.23% over eight test sets. The comparison results on varied prompts are shown in the Appendix Sec. A.
Table 3: Zero-shot FER results of our method with different instructions.
Static FER (UAR/WAR) Dynamic FER (UAR/WAR)
RAF-DB AffectNet-7 AffectNet-8 FERPlus DFEW FERV39k MAFW AFEW
ViT-B/32
Task-unrelated 33.70/34.77 30.08/30.09 26.10/26.11 33.03/35.43 21.57/20.74 19.02/17.69 14.71/17.51 29.42/30.88
Task-related 38.86/43.98 31.38/31.38 27.75/27.75 39.44/37.01 24.94/28.52 21.84/23.32 17.79/20.85 30.73/32.55
$\Uparrow$ 5.16/9.21 1.30/1.30 1.64/1.64 6.41/1.58 3.37/7.78 2.83/5.62 3.08/3.34 1.31/1.66
ViT-B/16
Task-unrelated 41.51/40.97 36.27/36.28 31.29/31.30 38.70/48.27 24.69/28.20 21.67/23.86 17.19/19.69 30.20/31.06
Task-related 48.84/54.68 39.46/39.46 34.04/34.04 40.19/53.82 26.40/31.64 25.11/28.57 18.98/23.64 31.50/32.63
$\Uparrow$ 7.33/13.71 3.19/3.18 2.74/2.74 1.49/5.55 1.71/3.44 3.44/4.71 1.79/3.96 1.30/1.58
ViT-L/14
Task-unrelated 57.32/61.26 43.14/43.14 37.20/37.20 46.79/54.16 37.94/43.40 25.86/30.23 23.54/26.47 38.08/40.33
Task-related 58.08/65.73 44.41/44.41 38.24/38.24 47.70/55.56 40.59/47.19 27.90/31.51 24.16/27.27 39.81/41.21
$\Uparrow$ 0.76/4.47 1.27/1.27 1.04/1.04 0.90/1.40 2.65/3.79 2.04/1.27 0.62/0.81 1.73/0.87
Evaluation of Different Instructions: To verify the effectiveness of the instruction adopted in our method, we conducted experiments on two types of instructions: task-related and task-unrelated. Regarding the task-unrelated instructions, we consider both empty and random text as input and take the average of them as the final results. For the task-related instruction, we provide results with three instructions and their average. As shown in Tab. 3, the task-related instructions are superior to task-unrelated instructions on both SFER and DFER, in which there are 3.14% average UAR improvements and 4.02% average WAR improvements over eight test sets with ViT-B-32, 2.87% average UAR improvements and 4.86% average WAR improvements over eight test sets with ViT-B-16, and 1.38% average UAR improvements and 1.87% average WAR improvements over eight test sets with ViT-L-14. We believe that task-related instructions facilitate LLMs to generate semantic features with task-aware ability, providing better objects for projection head optimisation. The detailed results on task-related and task-unrelated instructions are shown in the Appendix Sec. B.
Table 4: Zero-shot FER results of our method with different projection heads. 2-MLP and 3-MLP denote 2-layer and 3-layer multilayer perceptron, respectively. TE denotes Transformer encoder. LM denotes the learnable linear matrix.
Projection Heads Static FER (UAR/WAR) Dynamic FER (UAR/WAR)
RAF-DB AffectNet-7 AffectNet-8 FERPlus DFEW FERV39k MAFW AFEW
Evaluation of Different Projection Heads: In our method, to align the original CLIP feature representations into task-specific LLM knowledge space, we propose to utilize a projection head as a linear mapping to fulfil this. To investigate if the improvement of the complexity of the projectors can further improve the alignment, we also conduct experiments with 2-layer MLP, 3-layer MLP, and Transformer Encoders. As shown in Tab. 4, even with a simply learnable matrix and fewer parameters, our strategy is more powerful and efficient. Compared to the Transformer Encoders, except for the WAR results on FERV39k, ours shows much better results than more complex Transformer Encoders. We believe that more complex non-linear models will entail overfitting to the LLMs feature space, performing worse in downstream zero-shot FER tasks. The linear projection head adopted in our method is more suitable and interpretable for mapping the original CLIP feature space to a task-specific one, maintaining the capacity of the existing model to capture robust, versatile representations while tailoring it to the FER domain with minimal trainable parameters.
Evaluation of Varied Image and Text Encoders: To investigate if the LLM can do better than our method on zero-shot FER tasks, we conducted experiments using the LLM as either an image encoder (with I2T module), text encoder, or both. As shown in Tab. 5, adopting the LLM as either an image encoder or text encoder results in a worse zero-shot preference. Even when using the LLM as both an image and text encoder, the results remain undesirable and significantly lower than the plain CLIP model and our method. We believe the LLM excels in the semantic perception of images but performs poorly in recognition tasks. By contrast, based on the CLIP model, our method is more efficient and accurate in zero-shot FER tasks.
Table 5: Zero-shot FER results of varied image and text encoders. PH denotes the projection head trained in our method.
Image Encoder Text Encoder Static FER Dynamic FER
RAF-DB AffectNet-7 AffectNet-8 FERPlus DFEW FERV39k MAFW AFEW
4.4 Visualizations
To prove the benefit of the enhanced CLIP visual features, we utilize t-SNE [31] to illustrate the feature distribution of the original CLIP feature representations and projected feature representations. Different from the supervised methods learned with annotations which can gather samples of the same category compactly, zero-shot visual feature representations are learned by aligning the cross-modality representations, presenting a weak performance in gathering samples compactly. Despite this, from the visual feature distributions in Fig. 3, we can see the samples of the same category can be better gathered after the projection, especially for emotions of happiness and disgust. This demonstrates the model learns more discriminative latent visual features after the projection head, and can also explain why our method obtains superior results to the original plain CLIP model. In addition, we calculate the normalized feature variance of each emotion on FERPlus in the Appendix. The projected features demonstrate a lower variance than the CLIP feature space, proving the projected features are more concentrated. We also provide the confusion matrix comparison in the Appendix, which shows that our method captures more nuanced facial expression features.
5 Limitations
The proposed Exp-CLIP improves zero-shot FER results by incorporating domain knowledge from LLMs. However, our method still faces several limitations due to the following assumptions: 1) LLMs can adequately capture and describe the subtle nuances of facial expressions, and 2) LLMs can generate outputs that are well-grounded in the input data, minimizing hallucinations. To address 1), we utilize the Flan-T5-XXL model with 11B parameters, which provides more reliable and detailed knowledge. For 2), we use feature representations extracted after the LLM encoder, bypassing the decoder to mitigate the influence of hallucinations. In addition, our method captures more nuanced facial expression features derived from LLMs, e.g. performing better in distinguishing between neutral and other emotions. However, our zero-shot method still makes some incorrect predictions, particularly when distinguishing between similar emotions like Anger and Disgust, as well as Surprise and Fear, especially on hard or ambiguous samples.
6 Conclusion
Acknowledgments: This research utilised Queen Mary’s Apocrita HPC facility, supported by QMUL Research-IT. Zengqun Zhao is funded by Queen Mary Principal’s PhD Studentships. We thank Dr. Ziquan Liu, James Oldfield, and Yimeng Gu for their valuable comments.
A Comparison with CLIP on Varied Prompts